in triangle abc, the measure of angle b is 20 degrees more than three times the measure of angle a. the measure of angle c is 55 degrees more than the measure of angle a. find the measure of each angle.
Found 2 solutions by richwmiller, TheProdicalSon:
Answer by richwmiller(17219)

A triangle has 180 in its three interior angles
a+b+c=180
b=3a+20
c=55+a
substitute
a+3a+20+55+a=180
combine
5a+75=180
5a=105
a=21
b=3a+20=63+20=83
c=55+21=76
check
21+83+76=180
104+76=180
180=180
ok



Answer by TheProdicalSon(34)

b = 3a + 20
c = a + 55
a + b + c = 180
-c + a = -55
a + b + c = 180
-2a - b = -125
b - 3a = 20
-5a = -105
a = 21 degrees
b = 83 degrees
c = 76 degrees